OpenStudy (narissa):

what is square root 80 in simplified radical form

OpenStudy (narissa):

@GretaKnows

jimthompson5910 (jim_thompson5910):

hint: 80 = 16*5

OpenStudy (narissa):

is it 10 square root 8?

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{80} = \sqrt{16*5}\] \[\Large \sqrt{80} = \sqrt{16}*\sqrt{5}\] \[\Large \sqrt{80} = ???\]

OpenStudy (narissa):

16*5=80

jimthompson5910 (jim_thompson5910):

finish up my steps, what's missing?

OpenStudy (narissa):

165?

jimthompson5910 (jim_thompson5910):

what is the square root of 16?

OpenStudy (jdoe0001):

hint: \(\bf 80\to 2\cdot 2\cdot 2\cdot 2\cdot 5\to 2^2\cdot 2^2\cdot 5\)

OpenStudy (narissa):

8

jimthompson5910 (jim_thompson5910):

8 times 8 is not 16

jimthompson5910 (jim_thompson5910):

fill in the blank with the same number ____ times _____ = 16

OpenStudy (narissa):

4

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

so we get \[\Large \sqrt{80} = \sqrt{16*5}\] \[\Large \sqrt{80} = \sqrt{16}*\sqrt{5}\] \[\Large \sqrt{80} = 4*\sqrt{5}\]

OpenStudy (narissa):

so 4 and then square root 5

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (narissa):

thank u